How many different cash amounts can be made by choosing 4 coins from a penny, nickel, dime, quarter, loonie and a tonnie?
Would this be a combination?
I.E. 6C4

Answer by jim_thompson5910(35256)

Yes, then you evaluate that to get

n C r = (n!)/(r!(n-r)!)

6 C 4 = (6!)/(4!*(6-4)!)

6 C 4 = (6!)/(4!*2!)

6 C 4 = (6*5*4!)/(4!*2!)

6 C 4 = (6*5)/(2!)

6 C 4 = (6*5)/(2*1)

6 C 4 = (30)/(2)

6 C 4 = 15


So there are 15 combinations (order doesn't matter)